find the equation of the line that contains (-2, 8) and has a slope 1/2.
thanks a lot for helping me out! :)

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



The equation of a line containing a point and having a slope, m, is given by:




Just plug in your coordinates and you have your equation. You should probably manipulate the result so that your final answer is in either:

slope-intercept form: or

standard form:


John



Answer by stanbon(75887)

find the equation of the line that contains (-2, 8) and has a slope 1/2
------------------
m = 1/2, and y=8 when x = -2; so solve for b in the form y = mx + b
--------------
8 = (1/2)*-2 + b
b = 9
===========
Equation:
y = (1/2)x + 9
==================
Cheers,
Stan H.